OpenStudy (anonymous):

Let f(x)=(x−2)2 . Find the average rate of change of f(x) with respect to x from x=a to x=a+h if h≠0 . Simplify your answer as much as possible.

OpenStudy (anonymous):

Not getting the right answer again

OpenStudy (anonymous):

your job is to compute \[\frac{f(a+h)-f(a)}{h}\] rigth?

OpenStudy (anonymous):

except for the denominator i have a+h-(a)

OpenStudy (anonymous):

in this case \[f(x)=(x-2)^2=x^2-4x+4\] so compute \[\frac{(a+h)^2-4(a+h)+4-(a^2-4a+4)}{h}\] a bunch of algebra now

OpenStudy (anonymous):

I got -6a+h+4

OpenStudy (anonymous):

which is wrong

OpenStudy (anonymous):

it is clear than a + h - a = h rigtht?

OpenStudy (anonymous):

o yeah... nevermind then

OpenStudy (anonymous):

but yeah, i get how to do it. just getting the wrong answer for some reason.

OpenStudy (anonymous):

yes that is incorrect. you need to expand in the numerator and get \[\frac{a^2+2ah+h^2-4a-4h+4-a^2+4a-4}{h}=\frac{2ah-4h+h^2}{h}=2a-4+h\]

OpenStudy (anonymous):

why is the numerator (towards the end) PLUS 4a and MINUS 4 instead of the opposite thing (which is what I had).

OpenStudy (anonymous):

*?

OpenStudy (anonymous):

because it is \[-f(a)\] so you have to subtract the whole thing. in other words if \[f(a)=a^2-4a+4\] then \[-f(a)=-(a^2-4a+4)=-a^2+4a-4\]

OpenStudy (anonymous):

that always confuses me because i just think of it as a minus sign. i need to start putting everything in parenthesis. THANKS